# Adaptive RKHS Fourier Features for Compositional Gaussian Process Models

### v1

@@ Abstract
Deep Gaussian Processes (DGPs) leverage a compositional structure to model non-stationary processes. DGPs typically rely on local inducing point approximations across intermediate GP layers. Recent advances in DGP inference have shown that incorporating global Fourier features from Reproducing Kernel Hilbert Space (RKHS) can enhance the DGPs’ capability to capture complex non-stationary patterns. This paper extends the use of these features to compositional GPs involving linear transformations. In particular, we introduce Ordinary Differential Equation (ODE) -based RKHS Fourier features that allow for adaptive amplitude and phase modulation through convolution operations. This convolutional formulation relates our work to recently proposed deep latent force models, a multi-layer structure designed for modelling nonlinear dynamical systems. By embedding these adjustable RKHS Fourier features within a doubly stochastic variational inference framework, our model exhibits improved predictive performance across various regression tasks.
@@ 1 Introduction
Gaussian Processes (GPs) provide a principled Bayesian framework for function approximation, making them particularly useful in many applications requiring uncertainty calibration (Rasmussen and Williams, 2006), such as Bayesian optimisation (Snoek et al., 2012) and time-series analysis (Roberts et al., 2013). Despite offering reasonable uncertainty estimation, shallow GPs often struggle to model complex, non-stationary processes present in practical applications. To overcome this limitation, Deep Gaussian Processes (DGPs) employ a compositional architecture by stacking multiple GP layers, thereby enhancing representational power while preserving the model’s intrinsic capability to quantify uncertainty (Damianou and Lawrence, 2013). However, the conventional variational formulation of DGPs heavily depends on local inducing point approximations across intermediate GP layers (Titsias, 2009; Salimbeni and Deisenroth, 2017), which hinder the model from capturing the global structures commonly found in real-world scenarios.
Incorporating Fourier features into GP models has shown promise in addressing this challenge in GP inference due to the periodic nature of these features. A line of research uses Random Fourier Features (RFFs, (Rahimi and Recht, 2007)) of stationary kernels to convert the original (deep) GPs into Bayesian networks in weight space (Lázaro-Gredilla et al., 2010; Gal and Turner, 2015; Cutajar et al., 2017). Building on this concept within a sparse variational GP framework, recent advancements in inter-domain GPs (Lázaro-Gredilla and Figueiras-Vidal, 2009a; Van der Wilk et al., 2020) directly approximate the posterior of the original GPs by introducing fixed Variational Fourier Features (VFFs) through process projection onto a Reproducing Kernel Hilbert Space (RKHS)(Hensman et al., 2018; Rudner et al., 2020).
@@
In this paper, we focus on a type of GP characterised as the output of a convolution operation between a smoothing kernel and a latent GP. An example of this construction is the Latent Force Model (LFM, (Alvarez et al., 2009)), in which the smoothing kernel corresponds to the Green’s function associated with an Ordinary Differential Equation (ODE). By incorporating RKHS Fourier features into this framework, we derive adaptive global features inspired by the ODE, allowing for the optimisation of amplitudes and phases. We name the obtained features Variational Fourier Response Features (VFRFs) since they are derived from the output of a linear system. To enhance the capability of our model, we further use these adaptive features in a compositional GP model that stacks multiple LFMs, also known as Deep LFM (DLFM, (McDonald and Álvarez, 2021)). This hierarchical structure facilitates more precise and robust modelling of complex, non-stationary data. Our experimental results on both synthetic and real-world data demonstrate that incorporating these ODE-inspired RKHS Fourier features improves upon the standard practice of using VFFs.
@@ 2 Background
This section reviews concepts and preliminaries relevant to this work, including GPs, LFMs (Section  2.1), and the sparse variational framework with Fourier features for GP inference (Section 2.2). We also establish the notation used throughout the subsequent discussions.
2.1 Latent Force Models
Gaussian Processes
@@ 2.1 Latent Force Models
A Gaussian Process (GP) $f(\cdot)\sim\mathcal{GP}(m(\cdot),k(\cdot,\cdot^{\prime}))$ places probability measures on a function space $\{f:\mathbb{R}^{D}\rightarrow\mathbb{R}\}$ (Rasmussen and Williams, 2006). Its behaviour is characterised by the mean function $m:\mathbb{R}^{D}\rightarrow\mathbb{R}$ and the covariance function $k:\mathbb{R}^{D}\times\mathbb{R}^{D}\rightarrow\mathbb{R}$. The evaluation of the function $f(\cdot)$ at an input of interest $\mathbf{x}$ is a random variable denoted as $f(\mathbf{x})\in\mathbb{R}$. Given a dataset of inputs $\mathbf{X}=[\mathbf{x}_{n}]_{n=1}^{N}\in\mathbb{R}^{N\times D}$ and corresponding measurements $\mathbf{y}=[y_{n}]_{n=1}^{N}\in\mathbb{R}^{N}$, we assume $\mathbf{y}$ is observed from a noise-corrupted GP: $y_{n}=f(\mathbf{x}_{n})+\epsilon,\epsilon\sim\mathcal{N}(\epsilon\mid 0,\varepsilon^{2})$, where $\varepsilon^{2}$ is the noise variance. Then the posterior distribution $p(f\mid\mathbf{y})$ at any input locations $\mathbf{x},\mathbf{x}^{\prime}$ remains a GP $f\mid\mathbf{y}\sim\mathcal{GP}(\mu(\mathbf{x}),\Sigma(\mathbf{x},\mathbf{x}^{\prime}))$, with updated mean and covariance given by
$\mu(\mathbf{x})$ $=m(\mathbf{x})+k(\mathbf{x},\mathbf{X})(\mathbf{K}_{\mathbf{XX}}+\varepsilon^{2}\mathbb{I}_{N})^{-1}(\mathbf{y}-m(\mathbf{X})),$ (1)
$\Sigma(\mathbf{x},\mathbf{x}^{\prime})$ $=k(\mathbf{x},\mathbf{x}^{\prime})-k(\mathbf{x},\mathbf{X})(\mathbf{K}_{\mathbf{XX}}+\varepsilon^{2}\mathbb{I}_{N})^{-1}k(\mathbf{X},\mathbf{x}^{\prime}),$
where $[m(\mathbf{X})]_{n}=m(\mathbf{x}_{n})$, $[\mathbf{K_{XX}}]_{i,j}=k(\mathbf{x}_{i},\mathbf{x}_{j})$. The kernel parameters and the noise variance are typically optimised by maximising the log-likelihood $\log p(\mathbf{y}\mid\mathbf{X})$.
Latent Force Model
@@
A Latent Force Model (LFM, (Alvarez et al., 2009)) integrates differential equations with GPs to probabilistically model dynamic physical systems. Based on prior physical knowledge, a single-output LFM assumes the system’s output $f(t)$ is influenced by $Q$ latent forces $\{u_{q}(t)\}_{q=1}^{Q}$ through differential equations. Commonly, a first-order LFM uses the following form of ODE (Guarnizo and Álvarez, 2018)
@@
where $\gamma>0$ is a decay parameter, and $S_{q}\in\mathbb{R}$ is a sensitivity parameter. The solution for the output process $f(t)$ takes the form of weighted convolution integrals $f(t)=\sum_{q=1}^{Q}S_{q}\int_{0}^{t}G(t-\tau)u_{q}(\tau)\mathop{}\!\mathrm{d}\tau$, where $G(\cdot)$ denotes the Green’s function associated with the ODE.
Latent forces are presumed to follow GP priors, $u_{q}(t)\sim\mathcal{GP}(0,k_{q}(t,t^{\prime}))$, leading to a covariance function for the outputs $k_{f}(t,t^{\prime})=\sum_{q=1}^{Q}S^{2}_{q}\int_{0}^{t}\int_{0}^{t^{\prime}}G(t-\tau)k_{q}(\tau,\tau^{\prime})G(t^{\prime}-\tau^{\prime})\mathop{}\!\mathrm{d}\tau\mathop{}\!\mathrm{d}\tau^{\prime}.$ For some types of covariance functions $k_{q}(t,t^{\prime})$, e.g., the radial basis function (RBF), the covariance $k_{f}(t,t^{\prime})$ can either be computed explicitly (Lawrence et al., 2006) or approximated by the inner product using convolved Random Fourier Features (RFFs) (Rahimi and Recht, 2007)
$\boldsymbol{\phi}_{q}(t)=\frac{S_{q}}{\sqrt{M}}\left[\int_{0}^{t}G(t-\tau)e^{j\omega_{m}\tau}\mathop{}\!\mathrm{d}\tau\right]_{m=1}^{M},\quad\omega_{m}\overset{\mathrm{i.i.d.}}{\sim}p_{q}(\omega),$ (3)
where $j^{2}=-1$ and $\omega_{m}$ is sampled i.i.d. from the spectral density $p_{q}(\omega)$ associated with $k_{q}$ (Guarnizo and Álvarez, 2018). By plugging the physics-informed kernels into the GP posterior (1), LFMs embed domain-specific knowledge into the learning process and can utilise the sparse approximation techniques in GP inference.
2.2 Variational Fourier Features
Sparse Variational GPs
Exact GPs, as described in (1), suffer from $\mathcal{O}(N^{3})$ time complexity and are limited to Gaussian likelihoods. Sparse Variational Gaussian Processes (SVGPs, (Titsias, 2009; Hensman et al., 2013, 2015)) provide a scalable inference framework by introducing a small set of $M(\ll N)$ inducing points $\mathbf{Z}=[\mathbf{z}_{m}]_{m}^{M}\in\mathbb{R}^{M\times D}$ and corresponding inducing variables $\mathbf{u}=[u(\mathbf{z}_{m})]_{m=1}^{M}\in\mathbb{R}^{M}$ from the GP prior, i.e., $p(\mathbf{u})=\mathcal{N}(\mathbf{u}\mid m(\mathbf{Z}),\mathbf{K_{ZZ}})$. A variational distribution $q(\mathbf{u})=\mathcal{N}(\mathbf{u}\mid\mathbf{m},\mathbf{S})$ is employed to approximate the posterior process $q(f(\mathbf{x}))=\int p(f(\mathbf{x})\mid\mathbf{u})q(\mathbf{u})\mathop{}\!\mathrm{d}\mathbf{u}=\mathcal{N}(f\mid\tilde{m},\tilde{\Sigma})$, where
$\tilde{\mu}(\mathbf{x})$ $=m(\mathbf{x})+k(\mathbf{x},\mathbf{Z})\mathbf{K}^{-1}_{\mathbf{ZZ}}(\mathbf{m}-m(\mathbf{Z})),$ (4)
$\tilde{\Sigma}(\mathbf{x},\mathbf{x}^{\prime})$ $=k(\mathbf{x},\mathbf{x}^{\prime})+k(\mathbf{x},\mathbf{Z})\mathbf{K}^{-1}_{\mathbf{ZZ}}(\mathbf{S}-\mathbf{K_{ZZ}})\mathbf{K}^{-1}_{\mathbf{ZZ}}k(\mathbf{Z},\mathbf{x}^{\prime}).$
Variational Fourier Features
Inter-domain GPs (Lázaro-Gredilla and Figueiras-Vidal, 2009b; Álvarez et al., 2010; Van der Wilk et al., 2020) extend the domain of inducing variables by integrating the GP $f$ with a deterministic inducing function $g$:
This formulation allows for a redefinition of inducing variables $\mathbf{u}=u(\mathbf{Z})$, which still share the GP prior, albeit with alternative expressions of $k(\mathbf{x},\mathbf{Z})$ and $\mathbf{K_{ZZ}}$ used in (4). By choosing various functions for $g$, inter-domain GPs facilitate the construction of vector-valued basis functions $k(\cdot,\mathbf{Z})$ for more informative feature extraction while maintaining the standard SVGP framework.
VFFs (Hensman et al., 2018) define each inter-domain inducing variable $u_{m}$ of $\mathbf{u}$ by projecting the original GP $f$ onto a Fourier basis: $u_{m}=\langle\phi_{m},f\rangle_{\mathcal{H}}$, where $\langle\cdot,\cdot\rangle_{\mathcal{H}}$ denotes the Matérn RKHS inner product on an interval $[a,b]$, and $\phi_{m}$ is the $m$-th entry of a truncated Fourier basis
$\boldsymbol{\phi}(x)=[1,\cos(z_{1}(x-a)),\cdots,\cos(z_{M}(x-a)),\sin(z_{1}(x-a)),\cdots,\sin(z_{M}(x-a))],$ (6)
with $x$ a scalar input. In this setting, $\mathbf{z}=[z_{m}]_{m=1}^{M}$ are $M$ inducing frequencies, analogous to inducing points in the SVGP context  [^1] In Hensman et al. (2018), inducing frequencies are denoted as $\omega_{m}$. In this work, we use $z_{m}$ instead to maintain a notation consistent with the inducing points in SVGPs and inter-domain GPs.. The projection of $f$ onto this basis results in sinusoidal terms in the cross-covariance, i.e., $\mathrm{Cov}[f(\cdot),u_{m}]=\phi_{m}(\cdot)$ due to the reproducing property of the RKHS.
@@ 3 Methodology
This section describes the method of integrating RKHS Fourier features into compositional inter-domain GPs, with a specific focus on LFMs within SVGP framework. We start by adapting the conventional ODE used in LFMs to incorporate the inter-domain GPs with VFFs as a special instance of our model (Section 3.1). Details on VFRFs are provided in Section 3.2. We then extend our model from a single-layer to a hierarchical structure in Section 3.3.
@@ 3.1 LFMs with Modified ODEs
$\text{Cov}[f(t),f(t^{\prime})]=\int_{-\infty}^{t^{\prime}}\int_{-\infty}^{t}G(t-\tau)k(\tau,\tau^{\prime})G(t^{\prime}-\tau^{\prime})\mathop{}\!\mathrm{d}\tau\mathop{}\!\mathrm{d}\tau^{\prime}.$ (9)
The covariance function (also LFM kernel) can be calculated analytically if $k$ is a Matérn kernel with half-integer orders. We give the closed-form covariances in Table 1 of Appx. 9.
@@
The Green’s function $G(\cdot)$, determined by the system’s dynamics, serves as a signal filter. It effectively acts as a low-pass filter described by the ODE (7), with $\gamma$ representing the “cutoff frequency”. Mathematically, the convolution operator $G(\cdot)$ of the modified ODE will behave like the Dirac delta function in (8) as $\alpha=1,\beta\rightarrow 0^{+}$ (i.e.,$\gamma\rightarrow+\infty$), causing $f(t)$ to closely replicate $u(t)$. Fig. 1, left, illustrates this behaviour. We use the Matérn-$\tfrac{1}{2}$ kernel for the covariance $k(t,t^{\prime})$ of the latent force. The figure shows this covariance function (dashed) and two LFMs covariance functions (solid) $k_{f}(t,t^{\prime})$ with different $\gamma$ values. The LFM kernel reverts to the latent force kernel as $\gamma$ increases. However, the conventional LFM kernel without the ODE modification, i.e., (2) (Guarnizo and Álvarez, 2018) will get flattened since the corresponding Green’s function $\exp{(-\gamma t)}$ does not effectively mimic a valid Dirac delta function.
3.2 Variational Fourier Response Features for LFMs

### v2

@@ Abstract
Deep Gaussian Processes (DGPs) leverage a compositional structure to model non-stationary processes. DGPs typically rely on local inducing point approximations across intermediate GP layers. Recent advances in DGP inference have shown that incorporating global Fourier features from the Reproducing Kernel Hilbert Space (RKHS) can enhance the DGPs’ capability to capture complex non-stationary patterns. This paper extends the use of these features to compositional GPs involving linear transformations. In particular, we introduce Ordinary Differential Equation(ODE)–based RKHS Fourier features that allow for adaptive amplitude and phase modulation through convolution operations. This convolutional formulation relates our work to recently proposed deep latent force models, a multi-layer structure designed for modelling nonlinear dynamical systems. By embedding these adjustable RKHS Fourier features within a doubly stochastic variational inference framework, our model exhibits improved predictive performance across various regression tasks.
@@ 1 INTRODUCTION
Gaussian Processes (GPs) provide a principled Bayesian framework for function approximation, making them particularly useful in many applications requiring uncertainty calibration (Rasmussen and Williams,, 2006), such as Bayesian optimisation (Snoek et al.,, 2012) and time-series analysis (Roberts et al.,, 2013). Despite offering reasonable uncertainty estimation, shallow GPs often struggle to model complex, non-stationary processes present in practical applications. To overcome this limitation, Deep Gaussian Processes (DGPs) employ a compositional architecture by stacking multiple GP layers, thereby enhancing representational power while preserving the model’s intrinsic capability to quantify uncertainty (Damianou and Lawrence,, 2013). However, the conventional variational formulation of DGPs heavily depends on local inducing point approximations across GP layers (Titsias,, 2009; Salimbeni and Deisenroth,, 2017), which hinder the model from capturing the global structures commonly found in real-world scenarios.
Incorporating Fourier features into GP models has shown promise in addressing this challenge in GP inference due to the periodic nature of these features. A line of research uses Random Fourier Features (RFFs) (Rahimi and Recht,, 2007) of stationary kernels to convert the original (deep) GPs into Bayesian networks in weight space (Lázaro-Gredilla et al.,, 2010; Gal and Turner,, 2015; Cutajar et al.,, 2017). Building on this concept within a sparse variational GP framework, recent advancements in inter-domain GPs (Lázaro-Gredilla and Figueiras-Vidal, 2009a, ; Van der Wilk et al.,, 2020) directly approximate the posterior of the original GPs by introducing fixed Variational Fourier Features (VFFs) through process projection onto a Reproducing Kernel Hilbert Space (RKHS)(Hensman et al.,, 2018; Rudner et al.,, 2020).
@@
In this paper, we focus on a type of GP characterised as the output of a convolution operation between a smoothing kernel and a latent GP. An example of this construction is the Latent Force Model (LFM) (Alvarez et al.,, 2009), in which the smoothing kernel corresponds to the Green’s function associated with an Ordinary Differential Equation (ODE). By incorporating RKHS Fourier features into this framework, we derive adaptive global features inspired by the ODE, allowing for the optimisation of amplitudes and phases. We name the obtained features Variational Fourier Response Features (VFRFs) since they are derived from the output of a linear system. To enhance the capability of our model, we further use these adaptive features in a compositional GP model that stacks multiple LFMs, also known as Deep LFM (DLFM) (McDonald and Álvarez,, 2021). This hierarchical structure facilitates more precise and robust modelling of complex, non-stationary data. Our experimental results on both synthetic and real-world data demonstrate that incorporating these ODE-inspired RKHS Fourier features improves upon the standard practice of using VFFs.
@@ 2 BACKGROUND
This section reviews concepts and preliminaries relevant to this work and establishes the notation used throughout the subsequent discussions.
2.1 Sparse Variational Gaussian Process
@@
A GP $f(\cdot)\sim\mathcal{GP}(m(\cdot),k(\cdot,\cdot^{\prime}))$ places probability measures on a function space $\{f:\mathbb{R}^{D}\rightarrow\mathbb{R}\}$ (Rasmussen and Williams,, 2006). Its behaviour is characterised by the mean function $m:\mathbb{R}^{D}\rightarrow\mathbb{R}$ and the covariance function $k:\mathbb{R}^{D}\times\mathbb{R}^{D}\rightarrow\mathbb{R}$. The evaluation of the function $f(\cdot)$ at an input of interest $\mathbf{x}$ is a random variable denoted as $f(\mathbf{x})\in\mathbb{R}$. Given a dataset of inputs $\mathbf{X}=[\mathbf{x}_{n}]_{n=1}^{N}\in\mathbb{R}^{N\times D}$ and the corresponding measurements $\mathbf{y}=[y_{n}]_{n=1}^{N}\in\mathbb{R}^{N}$, we assume $\mathbf{y}$ is observed from a noise-corrupted GP: $y_{n}=f(\mathbf{x}_{n})+\epsilon,\epsilon\sim\mathcal{N}(\epsilon\mid 0,\varepsilon^{2})$, where $\varepsilon^{2}$ is the noise variance. The exact inference for the posterior distribution $p(f\mid\mathbf{y})$ suffers from $\mathcal{O}(N^{3})$ time complexity and is limited to Gaussian likelihoods.
Sparse Variational Gaussian Processes (SVGPs) (Titsias,, 2009; Hensman et al.,, 2013, 2015) provide a scalable inference framework by introducing a small set of $M(\ll N)$ inducing points $\mathbf{Z}=[\mathbf{z}_{m}]_{m}^{M}\in\mathbb{R}^{M\times D}$ and the corresponding inducing variables $\mathbf{u}=[u(\mathbf{z}_{m})]_{m=1}^{M}\in\mathbb{R}^{M}$ from the GP prior, i.e., $p(\mathbf{u})=\mathcal{N}(\mathbf{u}\mid\mathbf{0},\mathbf{K_{ZZ}})$. A variational distribution $q(\mathbf{u})=\mathcal{N}(\mathbf{u}\mid\mathbf{m},\mathbf{S})$ is employed to approximate the posterior process $q(f(\mathbf{x}))=\int p(f(\mathbf{x})\mid\mathbf{u})q(\mathbf{u})\mathop{}\!\mathrm{d}\mathbf{u}=\mathcal{N}(f\mid\tilde{m},\tilde{\Sigma})$, where
$\tilde{\mu}(\mathbf{x})$ $=m(\mathbf{x})+k_{\mathbf{x}\mathbf{Z}}\mathbf{K}^{-1}_{\mathbf{ZZ}}\mathbf{m},$ (1)
$\tilde{\Sigma}(\mathbf{x},\mathbf{x}^{\prime})$ $=k_{\mathbf{x}\mathbf{x}^{\prime}}+\mathbf{k}_{\mathbf{x}\mathbf{Z}}\mathbf{K}^{-1}_{\mathbf{ZZ}}(\mathbf{S}-\mathbf{K_{ZZ}})\mathbf{K}^{-1}_{\mathbf{ZZ}}\mathbf{k}_{\mathbf{Z}\mathbf{x}^{\prime}}.$
2.2 Variational Fourier Features
Inter-domain GPs (Lázaro-Gredilla and Figueiras-Vidal, 2009b, ; Álvarez et al.,, 2010; Van der Wilk et al.,, 2020) extend the domain of inducing variables by integrating the GP $f$ with a deterministic inducing function $g$:
This formulation allows for a redefinition of inducing variables $\mathbf{u}=u(\mathbf{Z})$, which still share the GP prior, albeit with alternative expressions of $\mathbf{k}_{\mathbf{x}\mathbf{Z}}$ and $\mathbf{K_{ZZ}}$ used in (1). By choosing various functions for $g$, inter-domain GPs facilitate the construction of vector-valued basis functions $k(\cdot,\mathbf{Z})$ for more informative feature extraction while maintaining the standard SVGP framework.
VFFs (Hensman et al.,, 2018) define each inter-domain inducing variable $u_{m}$ of $\mathbf{u}$ by projecting the original GP $f$ onto a Fourier basis: $u_{m}=\langle\phi_{m},f\rangle_{\mathcal{H}}$, where $\langle\cdot,\cdot\rangle_{\mathcal{H}}$ denotes the Matérn RKHS inner product on an interval $[a,b]$, and $\phi_{m}$ is the $m$-th entry of a truncated Fourier basis
$\boldsymbol{\phi}(x)=$ $[1,\cos(z_{1}(x-a)),\cdots,\cos(z_{M}(x-a)),$ (3)
$\sin(z_{1}(x-a)),\cdots,\sin(z_{M}(x-a))],$
with $x$ a scalar input. In this setting, $\mathbf{z}=[z_{m}]_{m=1}^{M}$ are $M$ inducing frequencies, analogous to inducing points in the SVGP context. The projection of $f$ onto this basis results in sinusoidal terms in the cross-covariance, i.e., $\mathrm{Cov}[f(\cdot),u_{m}]=\phi_{m}(\cdot)$ due to the reproducing property of the RKHS.
2.3 Latent Force Model
@@
An LFM (Alvarez et al.,, 2009) is a GP model integrating differential equations to model dynamic physical systems probabilistically. Based on prior physical knowledge, a single-output LFM assumes the system’s output $f(t)$ is influenced by $Q$ latent forces $\{u_{q}(t)\}_{q=1}^{Q}$ through differential equations. Commonly, a first-order LFM uses the following form of ODE (Guarnizo and Álvarez,, 2018)
@@
where $\gamma>0$ is a decay parameter, and $S_{q}\in\mathbb{R}$ is a sensitivity parameter. The solution for the output $f(t)$ takes the form of weighted convolution integrals $f(t)=\sum_{q=1}^{Q}S_{q}\int_{0}^{t}G(t-\tau)u_{q}(\tau)\mathop{}\!\mathrm{d}\tau$, where $G(\cdot)$ denotes the Green’s function associated with the ODE.
Latent forces are presumed to follow GP priors, $u_{q}(t)\sim\mathcal{GP}(0,k_{q}(t,t^{\prime}))$, leading to a covariance function for the outputs $k_{f}(t,t^{\prime})=\sum_{q=1}^{Q}S^{2}_{q}\iint G(t-\tau)k_{q}(\tau,\tau^{\prime})G(t^{\prime}-\tau^{\prime})\mathop{}\!\mathrm{d}\tau\mathop{}\!\mathrm{d}\tau^{\prime}.$ For some types of covariance functions $k_{q}(t,t^{\prime})$, e.g., the radial basis function (RBF), $k_{f}(t,t^{\prime})$ can either be computed explicitly (Lawrence et al.,, 2006) or approximated by using convolved RFFs (Guarnizo and Álvarez,, 2018; Rahimi and Recht,, 2007). By plugging the physics-informed kernels into the GP posterior, LFMs embed domain-specific knowledge into the learning process and can utilise the sparse approximation techniques in GP inference.
@@ 3 METHODOLOGY
This section describes integrating RKHS Fourier features into compositional GPs, with a specific focus on LFMs within the SVGP framework. We start by adapting the conventional ODE used in LFMs to incorporate VFFs as a special instance of our model (Section 3.1). Details on VFRFs are provided in Section 3.2. We then extend our model from a single-layer to a hierarchical structure in Section 3.3.
@@ 3.1 LFMs with Modified ODEs
$\text{Cov}[f(t),f(t^{\prime})]=$ (7)
$\quad\int_{-\infty}^{t^{\prime}}\int_{-\infty}^{t}G(t-\tau)k(\tau,\tau^{\prime})G(t^{\prime}-\tau^{\prime})\mathop{}\!\mathrm{d}\tau\mathop{}\!\mathrm{d}\tau^{\prime}.$
The covariance function (also called LFM kernel in this paper) can be calculated analytically if $k$ is a Matérn kernel with half-integer orders. We give the closed-form covariances in Table 3 of Appendix C.
$t_{1}$$t_{2}$$f^{1}_{1}$$f^{1}_{2}$$\Sigma$$g^{1}_{1}$$g^{1}_{2}$$f^{2}_{1}$$f^{2}_{2}$$\Sigma$$g^{2}_{1}$$y$
(a) Inter-domain deep Gaussian process (IDDGP)
[…]
(b) Deep latent force model leveraging RFFs (DLFM-RFF)
[…]
(c) Our deep latent force model with VFRF (DLFM-VFRF)
Figure 2: A conceptual illustration of how our model (2(c)) differs from the IDDGP (2(a)) and the DLFM-RFF (2(b)). Compared to (2(a)), our model additionally applies convolution operators $G$ from the ODEs to each input dimension: $f(t)=\int G(t-\tau)u(\tau)\mathop{}\!\mathrm{d}\tau$, where $G(\cdot)$ represents the Green’s function and $u(\cdot)$ is a GP prior with Matérn kernels. Compared to (2(b)) using RFFs $\varphi(\cdot)$ for low-rank covariance matrix approximation and making inference over weights $W$, our model uses Fourier features derived from applying linear transformations to GPs and make inference in an inter-domain way. For a high-level comparison with other models, see Fig. 7.
@@
The Green’s function $G(\cdot)$, determined by the system’s dynamics, serves as a signal filter. It effectively acts as a low-pass filter described by the ODE (5), with $\gamma$ representing the “cutoff frequency”. Mathematically, the convolution operator $G(\cdot)$ of the modified ODE will behave like the Dirac delta function in (6) as $\alpha=1,\beta\rightarrow 0^{+}$ (i.e.,$\gamma\rightarrow+\infty$), causing $f(t)$ to closely replicate $u(t)$. Fig. 1, left, illustrates this behaviour. We use the Matérn-$\tfrac{1}{2}$ kernel for the covariance $k(t,t^{\prime})$ of the latent force. The figure shows this covariance function (dashed) and two LFMs covariance functions (solid) $k_{f}(t,t^{\prime})$ with different $\gamma$ values. The LFM kernel reverts to the latent force kernel as $\gamma$ increases. However, the conventional LFM kernel without the ODE modification, i.e., (4) (Guarnizo and Álvarez,, 2018) will get flattened since the corresponding Green’s function $\exp{(-\gamma t)}$ does not effectively mimic a valid Dirac delta function.
3.2 Variational Fourier Response Features